# Multi-Scale Spatial-Temporal Hypergraph Network with Lead-Lag Structures for Stock Time Series Forecasting

### v1

@@ B.3 Hypergraph-based Stock Time Series Forecasting
Traditional graphs have limitations when it comes to modeling higher-order multivariate relationships, as they can only represent relations between pairs of nodes. To address this issue, Hypergraph Neural Networks (HGNNs) (Feng et al., 2019c) were introduced. Hypergraphs extend the concept of simple graphs to enable the capture of relationships among multiple stocks (Chen et al., 2020). Hypergraphs have gained increasing attention and application across various fields. Hyperedge in a hypergraph represents a set of vertices, making them suitable for modeling non-pairwise relationships (Luo et al., 2014; Sawhney et al., 2020). For example, HGAM (Li et al., 2022) is a hypergraph-based reinforcement learning method for stock portfolio selection. THGNN (Xiang et al., 2022) is a temporal and heterogeneous graph neural network model that aims to learn dynamic relationships by using two-stage attention mechanisms. STHAN-SR (Sawhney et al., 2021) enhances corporate relevance based on Wiki data and uses hypergraph convolution to propagate information from higher-order neighbors. DHSTN (Liao et al., 2024) is a dynamic hypergraph network for learning spatio-temporal relations of stocks. Dynamic hypergraphs are generated adaptively using (Graph Attention Networks) GATs to capture time-varying higher-order spatial relationships among stocks. The latest method, ESTIMATE (Huynh et al., 2023), employs hypergraphs to capture non-pairwise correlations, utilizing temporal generative filters to identify individual patterns for each stock.
@@ B.5 Hyperparameter Sensitivity
This experiment addresses question RQ3 on the hyperparameter sensitivity. Due to space limitations, we focus on the most important hyperparameters—see Figure 5.
@@
Score weight $\alpha$: Besides, we adopt the score weight in section 3.7 to trade off the MSE loss and the ranking-aware loss—see Figure  5(d). Experimental results show that all datasets achieve optimal performance and exhibit high stability when the score weight is set to 1. Therefore, we recommend setting this parameter to 1 in future research to obtain more optimized results.

### v2

@@ B.3 Hypergraph-based Stock Time Series Forecasting
Traditional graphs have limitations when it comes to modeling higher-order multivariate relationships, as they can only represent relations between pairs of nodes. To address this issue, Hypergraph Neural Networks (HGNNs) Feng et al. [2019c] were introduced. Hypergraphs extend the concept of simple graphs to enable the capture of relationships among multiple stocks Chen et al. [2020]. Hypergraphs have gained increasing attention and application across various fields. Hyperedge in a hypergraph represents a set of vertices, making them suitable for modeling non-pairwise relationships Luo et al. [2014]; Sawhney et al. [2020]. For example, HGAM Li et al. [2022] is a hypergraph-based reinforcement learning method for stock portfolio selection. THGNN Xiang et al. [2022] is a temporal and heterogeneous graph neural network model that aims to learn dynamic relationships by using two-stage attention mechanisms. STHAN-SR Sawhney et al. [2021] enhances corporate relevance based on Wiki data and uses hypergraph convolution to propagate information from higher-order neighbors. DHSTN Liao et al. [2024] is a dynamic hypergraph network for learning spatio-temporal relations of stocks. Dynamic hypergraphs are generated adaptively using (Graph Attention Networks) GATs to capture time-varying higher-order spatial relationships among stocks. The latest method, ESTIMATE Huynh et al. [2023], employs hypergraphs to capture non-pairwise correlations, utilizing temporal generative filters to identify individual patterns for each stock.
@@ B.4 Illustration Figures
Figure 5 illustrates the specific lead-lag interaction process within the window, with a lead-lag step of 3 as an example. Figure 5 depicts three types of message passing: hypernodes for stocks and hyperedges for industries.
@@ B.5 Hyperparameter Sensitivity
We also study the parameter sensitivity of the Hermes—see Figure 3.
@@
Score weight $\alpha$: Besides, we adopt the score weight in section 3.7 to trade off the MSE loss and the ranking-aware loss—see Figure  3(d). Experimental results show that all datasets achieve optimal performance and exhibit high stability when the score weight is set to 1. Therefore, we recommend setting this parameter to 1 in future research to obtain more optimized results.